OpenStudy (anonymous):

Which value of x makes the following equation true? 3(x - 5) - 14 = 2(x - 7) + 3 A. 18 B. 25 C. 13 D. 11

OpenStudy (math&ing001):

You know how to solve for x ?

OpenStudy (math&ing001):

@PurplePanda312

OpenStudy (anonymous):

Sorta

OpenStudy (math&ing001):

Try and I'll help through.

OpenStudy (anonymous):

I don't know how to start

OpenStudy (math&ing001):

Ok, first use FOIL on 3(x-5) and 2(x-7). Tell me what you get.

OpenStudy (anonymous):

Would I add 3 and 5

OpenStudy (math&ing001):

Nope, this is FOIL: a(b+c) = ab + ac So 3(x-5) = 3x - 15 and 2(x-7) = 2x - 14 The equation becomes : 3x - 15 - 14 = 2x - 14 + 3 3x - 2x = 18 x = 18